A cyclist traveled at a rate of 24 mi/h to visit a nearby town. The cyclist averaged 18 mi/h on the return trip. If the round trip took 6 hours, find the one way distance to the nearby town.
Answer by checkley79(3341)

D=RT
D=24T THE FIRST TRIP.
D=18(6-T) THE RETURN TRIP.
24T=18(6-T)
24T=108-18T
24T+18T=108
42T=108
T=108/42
T=2.57 HOURS FOR THE FIRST TRIP.
6-2.57=3.43 HOURS FOR THE RETURN TRIP.
PROOF:
24*2.57=18*3.43
61.7=61.7 MILES ONE WAY.